How can I solve 1/4X with a exponent of -2?
Answer by Boreal(15235)

If this is (1/4x)^-2
it is 1^(-2)/(4x)^(-2)
That is (4x)^2=16x^2, because we flip a fraction that is raised to a negative exponent, and the exponent then becomes positive.
----------------------
If this is (1/4)x^(-2)=(x/4)^(-2)=(4/x)^2=16/(x^2)